OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

Here's my answer: Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. We will need to decide how much they are selling each flower for. It says they are charging what they paid, which is 50 cents per flower plus 40 percent. .5 + .4(.5) = 0.7 They are charging 70 cents per flower. How many flowers do they need to sell to get $500? 500 = x * .70 .7x = 500. X = 714.285714285714 (round to 715) They will need to sell 715 flowers.

ganeshie8 (ganeshie8):

partially correct

OpenStudy (anonymous):

Patially? what is the incorrect part?

ganeshie8 (ganeshie8):

` Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500.`

ganeshie8 (ganeshie8):

They want a profit of $500, not just sales

ganeshie8 (ganeshie8):

use below to setup the equation : profit = (selling price) - (cost price)

OpenStudy (anonymous):

500 = x - 50??

ganeshie8 (ganeshie8):

profit = (selling price) - (cost price) 500 = (x*0.7) - (x*0.5)

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

okay, one minute

OpenStudy (anonymous):

500 = (x⋅0.7)− (x⋅0.5) x = 2500

ganeshie8 (ganeshie8):

Excellent ! so they need to sell 2500 flowers to make a profit of 500

OpenStudy (anonymous):

i was sort of guessing though, i am an idiot.. :( idk what the steps are

OpenStudy (anonymous):

i mean, like.. i know theres pemdas but, i couldnt understand how to solve this.

OpenStudy (anonymous):

x * 0.7 = 0.7x x * 0.5 = 0.5x 0.7x - 0.5x 0.2x Right??

ganeshie8 (ganeshie8):

Looks perfect !

OpenStudy (anonymous):

okay thanks! but one more thing..

OpenStudy (anonymous):

this does not look like a full equation to me, i feel like im missing something

OpenStudy (anonymous):

division never came to my head im so dumb, lol

OpenStudy (anonymous):

thank you, i got it now.. 5 + .4(.5) = 0.7 profit = (selling price) - (cost price) x * 0.7 = 0.7x x * 0.5 = 0.5x 0.7x - 0.5x 0.2x 500 / 0.2 = 2500 X = 2500